# Leveraging Audio and Text Modalities in Mental Health: A Study of LLMs Performance

### v1

@@ 4.5.1 Model Exclusions and Limitations
In evaluating models for the study, each option was scrutinized based on its ability to analyze extended audio inputs and perform complex, context-heavy textual analysis. The research primarily focused on models capable of handling long audio files representing entire spoken paragraphs and deriving insights from complex conversational dynamics. Below is an overview of why certain models were excluded, highlighting their limitations in relation to the study’s requirements:
  • Qwen/Qwen2: This model was excluded due to a technical limitation related to the size of audio files it can process. Specifically, Qwen2-Audio, which supports an audio file size of up to 10240 KB. However, the dataset used in the study contained longer audio files representing entire spoken paragraphs, which exceeded this maximum size limit. This limitation made Qwen/Qwen2 unsuitable for the research, which required processing extended audio inputs to analyze spoken content effectively.
  • Flamingo: Although Flamingo is a powerful multimodal model that excels in combining image and text processing, it was excluded because its audio capabilities were not robust enough for the study’s focus. The research aimed to evaluate models designed for handling audio and text data, while Flamingo is more oriented toward tasks involving visual and textual data. Its strength lies in few-shot learning and integrating visual-textual data rather than deep audio processing, which made it less suitable for a comparative analysis of audio-based models.
@@ 4.6 Prompt Templates
This section details the specific prompts employed to direct the analysis by Large Language Models in our study. These prompts are crafted to instruct the LLMs on processing the provided inputs—whether text transcriptions or raw audio files—from participants. The primary objective of employing these prompts is to assess the presence of depression or PTSD symptoms accurately. Each prompt is designed to maximize the LLMs’ capabilities in interpreting and diagnosing based on the modality being tested. Through the systematic application of these prompts, we aim to not only quantify the performance differences between text and audio modalities but also to explore whether a multimodal approach can enhance the predictive accuracy and reliability of mental health assessments. This approach allows for a structured evaluation of how different data presentations can influence the performances of LLMs.
@@ 4.6.1 Binary Prompts
{mdframed}
[backgroundcolor=blue!10,linecolor=blue!10] PHQ_Binary (Depression) and PCL-C (PTSD) Prompts Prompt 1 (P1) The following (input) is from an interaction between an interviewer and an interviewee. These interviews were collected as part of a larger effort to create a computer agent that interviews people and identifies verbal and nonverbal indicators of mental illnesses. Guidelines: - Concise Response: Respond with one of these 2 words only ("Yes", or "No"). - ’Yes’ refers to the interviewee exhibiting (illness), and ’No’ refers that he has no clear symptoms. - No Explanations: Don’t provide explanations for your assessment. - Ambiguity: If the post is unclear, choose the most probable label. (input type of the interview) "(input)"
Prompt 2 (P2) Analyze the following interview (input) to determine if the interviewee exhibits symptoms of depression. Focus on language patterns, emotional expressions, and overall demeanor. Respond with ’Yes’ if (illness) is indicated, otherwise respond with ’No’. Respond with one word only! "(input)"
Prompt 3 (P3) As a psychologist, analyze the provided interview (text/audio): (input) Does the interviewee exhibit clear symptoms of (illness)? Answer with ’Yes’ or ’No’ only without explaining your reasoning; use only one word!
@@ 4.6.2 Severity Prompts
In the following section, figure 5 presents the PHQ_Score Mapped (Depression severity) Prompts, which are used to assess the severity of depression in participants. These prompts ask the model to categorize the level of depression based on the input provided, whether in text or audio format, using the PHQ-8 scoring system. The model classifies the participant’s depression into one of five categories: minimal or no depression (0), mild depression (1), moderate depression (2), moderately severe depression (3), and severe depression (4). These categories provide a nuanced view of the participant’s mental health by determining the extent of their depressive symptoms. In figures 5 6, the term "input" refers to either a transcription of the participant’s spoken responses or the original audio files used in the analysis.
{mdframed}
[backgroundcolor=blue!10,linecolor=blue!10] PHQ_Score Mapped (Depression Severity) Prompts Prompt 1 (P1) The following (input) is from an interaction between an interviewer and an interviewee. These interviews were collected as part of a larger effort to create a computer agent that interviews people and identifies verbal and nonverbal indicators of mental illnesses. The (input) is: ’{input}’. Based on the (input), classify the severity of depression into: 0 (Minimal or no depression): No or very few signs. 1 (Mild depression): Mostly mild symptoms. 2 (Moderate depression): Moderate symptoms impacting daily life. 3 (Moderately severe depression): Noticeable symptoms hindering daily life. 4 (Severe depression): Severe, pervasive symptoms greatly impairing function. Answer with one number only!
Prompt 2 (P2) Analyze the following interview (input) for symptoms of depression. Focus on language patterns, emotional expressions, and demeanor. The (input) is: ’{input}’. Classify severity into: 0 (Minimal or no depression): No or very few signs. 1 (Mild depression): Mostly mild symptoms. 2 (Moderate depression): Moderate symptoms impacting daily life. 3 (Moderately severe depression): Noticeable symptoms hindering daily life. 4 (Severe depression): Severe, pervasive symptoms greatly impairing function. Answer with one number only!
@@
{mdframed}
[backgroundcolor=blue!10,linecolor=blue!10] PTSD Severity Mapped Prompts Prompt 1 (P1) The following (input) is from an interaction between an interviewer and an interviewee. These interviews were collected as part of a larger effort to create a computer agent that interviews people and identifies verbal and nonverbal indicators of mental illnesses. The (input) is: ’{input}’. Based on the (input), classify the severity of PTSD into: 0 (Little to no severity): No significant signs of PTSD symptoms; little to no impact on daily life. 1 (Moderately severe): Moderate symptoms of PTSD with a noticeable impact on behavior and daily routines. 2 (High severity): Severe PTSD symptoms with significant impact, disrupting daily life and possibly requiring intervention. Answer with one number only, 0 to 2, corresponding to these categories.
Prompt 2 (P2) Analyze the following interview (input) to determine if the interviewee exhibits symptoms of PTSD. Focus on language patterns, emotional expressions, and overall demeanor. The (input) is: ’{input}’. Based on the (input), classify the severity of PTSD into: 0 (Little to no severity): No significant signs of PTSD symptoms; little to no impact on daily life. 1 (Moderately severe): Moderate symptoms of PTSD with a noticeable impact on behavior and daily routines. 2 (High severity): Severe PTSD symptoms with significant impact, disrupting daily life and possibly requiring intervention. Answer with one number only, 0 to 2, corresponding to these categories.
Figure 6: PTSD severity classification prompts
@@ 4.6.3 Multiclass Prompts
{mdframed}
[backgroundcolor=blue!10,linecolor=blue!10] Multiclass Prompts Prompt 1 (P1) The following (input) is from an interaction between an interviewer and an interviewee. These interviews were collected as part of a larger effort to create a computer agent that interviews people and identifies verbal and nonverbal indicators of mental illnesses. The (input) is: ’{input}’. The interviewee could have multiple illnesses at the same time. Analyze the transcript and identify if the interviewee has Depression, PTSD, both, or is Normal. Respond with only one of these words: "Depressed", "PTSD", "Normal", "Depressed and PTSD". Prompt 2 (P2) Analyze the following interview (input) to determine if the interviewee exhibits clear symptoms of Depression or PTSD according to the provided guidelines. Guidelines: The interviewee may have multiple illnesses at the same time or be normal. Concise Response: Respond with one of these four words only: "Depressed", "PTSD", "Depressed and PTSD", "Normal". No Explanations: Do not provide explanations for your assessment. Ambiguity: If the post is unclear, choose the most probable label. The (input) of the interview: ’{input}’. Only answer with one of the specified words.
@@ 4.6.4 Few-Shot Prompts
  • Severity and Multi-label Tasks: we strategically selected samples that were near-miss classifications by the model during zero-shot evaluations. This selection was based on samples where the model’s predictions were off by just one label, indicating a subtle misunderstanding of the distinguishing features between closely related classes. This methodological choice aimed to challenge the models with difficult examples where minor nuances in features are decisive, thereby enhancing the robustness and precision of the models through few-shot learning.
@@
{mdframed}
[backgroundcolor=blue!10,linecolor=blue!10] Example Few-shot Prompt (Prompt 1 for Binary Depression Detection) The following transcript is from an interaction between an interviewer and an interviewee. These interviews were collected as part of a larger effort to create a computer agent that interviews people and identifies verbal and nonverbal indicators of mental illnesses. If the text appears to be for a person who has Depression, answer with ’Yes’; if not, answer with ’No’. Only answer with Yes or No; respond with one word only! Here are 3 samples from these interviews and their labels. Use them as a reference: First sample transcription: (sample transcription) First sample label: No Second sample transcription: (sample transcription) Second sample label: Yes Third sample transcription: (sample transcription) Third sample label: Yes Label the following transcription: ’sample to be labeled’.
@@ 4.7 Parsers
Among the models used in this study, the larger models like Llama 3 70Band Phi-3.5-MoE performed well when following task guidelines, typically adhering to the expected response formats. Conversely, some smaller models, such as Gemma 2 9B and Phi-3.5-Mini, struggled to maintain this level of compliance. These models frequently provided additional explanations, deviating from the expected outputs. Smaller models were less consistent, requiring additional handling to extract the necessary information.
@@
  • Binary Detection: The parser was designed to search the LLM outputs for the text "yes" or "no." If one of these options appeared in the response, it was taken as the final output. If both "yes" and "no" appeared simultaneously, or if neither were found, the model output was deemed invalid.
  • Severity Detection: For this task, the parser focused on extracting numerical values that corresponded to the specified severity range. If a single valid number was detected, it was accepted as the answer. However, if multiple numbers appeared, or if the number fell outside of the allowed range, the output was flagged as invalid.
@@ 4.8 Evaluation Metrics
To evaluate the performance of the Large Language Models (LLMs) in tasks with uneven class distributions, as detailed in the dataset section (see Section 3), we prioritize Balanced Accuracy (BA) as our main evaluation metric. This measure, calculated as the average recall obtained across each class, fairly reflects the model’s effectiveness in identifying both prevalent and rare conditions. By employing Balanced Accuracy, we ensure a comprehensive evaluation of the LLMs’ capabilities.
@@
Here, $i$ represents an index for each class, ranging from 0 to $N-1$, where $N$ is the total number of classes. The recall for each class is computed to assess the model’s ability to correctly identify samples of that particular class.
@@
For binary classification tasks, we additionally use the F1 Score to evaluate model performance. The F1 Score is crucial as it provides a balance between precision and recall, making it a valuable metric for situations where the cost of false positives and false negatives is high. This metric is especially important in mental health assessments, where accurately distinguishing between conditions such as depression and PTSD is critical.
@@
While weighted F1 score is defined as
[…]
$w_{i}=\frac{\text{No. of samples in class }i}{\text{Total number of samples}}$
@@
To comprehensively assess the performance differences between each modality in our analysis, we employed two key metrics: the Modal Superiority Score (MSS) and the Disagreement Resolvement Score (DRS). These metrics are instrumental in quantifying the relative efficacy of individual and combined modalities in making correct predictions, particularly in the face of disagreement, by applying these metrics, which are detailed in the equations below, we gain valuable insights into which modalities perform better or worse and explore whether combining modalities enhances the predictive accuracy and giving insights on how a modality improves upon another.
Modal Superiority Score (MSS): The Modal Superiority Score (MSS) quantifies the net superiority of one modality over another by comparing how often each modality correctly predicts outcomes when the other does not. This metric can be applied not only to comparisons between individual modalities, such as Audio versus Text but also in evaluating the performance of a combined modality (e.g., Audio+Text) against individual modalities and the collective agreement of these modalities (e.g., cases where both Audio and Text are either correct or incorrect). This comprehensive application of MSS allows for assessing the relative strength of combined modalities over both their constituent individual modalities and their collective concordance. MSS values can be positive or negative, with a positive value indicating that modality A performs better than modality B, and a negative value suggesting the opposite.
$\text{MSS}_{A\text{vs}B}=\left(\frac{\text{Correctly predicted by }A\text{ andincorrectly by }B-\text{Correctly predicted by }B\text{ and incorrectly by }A}{\text{Total number of disagreements}}\right)\times 100\%$
@@
DRS=(Correctly ResolvedIncorrectly ResolvedTotal Number of $\text{DRS}=\left(\frac{\text{Correctly Resolved}-\text{Incorrectly Resolved}}{\text{Total Number of Disagreements}}\right)\times 100\%$
@@ 5 Results & Discussion
In the upcoming sections, we analyze the performances of various models for several mental health classification tasks, including Binary Depression, Binary PTSD, Depression Severity, and PTSD severity. The aim is to determine which models perform best under different conditions across both text and audio modalities, with no preprocessing or fine-tuning applied, relying solely on zero-shot inference.
@@ 5.1 Binary Depression Classification Results
Table 6: Report on model performances for Depression binary classification across text and audio modalities and their combination. Details about the specific prompts used can be found in Figure 4. The underlined bold values are the best score for the specific prompt.
@@
Text Model BA F1 BA F1 BA F1
Llama 3 70B 71.9% 0.61 74.7% 0.64 65.7% 0.54
@@
Audio Gemini 1.5 Pro 76.1% 0.66 74.5% 0.64 74% 0.64
@@
Audio and Text Gemini 1.5 Pro 74% 0.64 77% 0.67 77.3% 0.67
@@
The results presented in the Table 6 underscore the strong performances of the Gemini 1.5 Flash and Gemini 1.5 Pro across various modalities. Remarkably, the Gemini 1.5 Flash achieved the highest balanced accuracy of 77.4% and an F1 score of 0.68 in the combined modalities on Prompt 3, surpassing all other models. The Gemini 1.5 Pro also demonstrated exceptional capabilities, especially in the text modality, where it reached a balanced accuracy of 76.2% and an F1 score of 0.66 on Prompt 1, and comparable results in the audio modality.
In the text modality, the Gemini 1.5 Pro emerged as the top performer. GPT-4o mini also exhibited robust performance, achieving a balanced accuracy of 75.8% on Prompt 1 and an impressive F1 score of 0.64 on Prompt 3. Additionally, Phi-3.5-MoE proved highly effective, particularly with a balanced accuracy of 75.2% and an F1 score of 0.65 on Prompt 2. These models, including Gemini 1.5 Flash, consistently ranked high in text-based depression classification, highlighting their efficacy.
Overall, the combined use of text and audio modalities proved more effective, with Gemini 1.5 Flash leading the performance in these integrated assessments. Both modalities showed high efficacy in detecting depression, yet the multimodal approach allowed Gemini 1.5 Flash to achieve the highest scores, illustrating the advantage of leveraging both text and audio inputs. This demonstrates the robust capabilities of multimodal models in handling complex diagnostic tasks.

### v2

@@ 4.5.1 Model Exclusions and Limitations
In evaluating models for the study, each option was scrutinized based on its ability to analyze extended audio inputs and perform complex, context-heavy textual analysis. The research primarily focused on models capable of handling long audio files representing entire spoken paragraphs and deriving insights from complex conversational dynamics. Below is an overview of why specific models were excluded, highlighting their limitations about the study’s requirements:
  • Qwen/Qwen2: This model was excluded due to a technical limitation related to the size of audio files it can process. Specifically, Qwen2-Audio supports an audio file size of up to 10240 KB. However, the dataset used in the study contained longer audio files representing entire spoken paragraphs, which exceeded this maximum size limit. This limitation made Qwen/Qwen2 unsuitable for the research, which required processing extended audio inputs to analyze spoken content effectively.
  • Flamingo: Although Flamingo is a powerful multimodal model that excels in combining image and text processing, it was excluded because its audio capabilities were not robust enough for the study’s focus. The research aimed to evaluate models designed for handling audio and text data, while Flamingo is more oriented toward tasks involving visual and textual data. Its strength lies in few-shot learning and integrating visual-textual data rather than deep audio processing, which makes it less suitable for a comparative analysis of audio-based models.
@@ 4.6 Prompt Templates
This section details the specific prompts employed to direct the analysis by Large Language Models in our study. These prompts are crafted to instruct the LLMs on processing the provided inputs—whether text transcriptions or raw audio files—from participants. The primary objective of employing these prompts is to assess the presence of depression or PTSD symptoms accurately. Each prompt is designed to maximize the LLMs’ capabilities in interpreting and diagnosing based on the modality being tested. Through the systematic application of these prompts, we aim to not only quantify the performance differences between text and audio modalities but also to explore whether a multimodal approach can enhance the predictive accuracy and reliability of mental health assessments. This approach allows for a structured evaluation of how different data presentations can influence the performance of LLMs. ’
@@ 4.6.1 Binary Prompts
Prompt 1 (P1) The following (input) is from an interaction between an interviewer and an interviewee. These interviews were collected as part of a larger effort to create a computer agent that interviews people and identifies verbal and nonverbal indicators of mental illnesses. Guidelines: - Concise Response: Respond with one of these 2 words only ("Yes", or "No"). - ’Yes’ refers to the interviewee exhibiting (illness), and ’No’ refers that he has no clear symptoms. - No Explanations: Don’t provide explanations for your assessment. - Ambiguity: If the post is unclear, choose the most probable label. (input type of the interview) "(input)" Prompt 2 (P2) Analyze the following interview (input) to determine if the interviewee exhibits symptoms of depression. Focus on language patterns, emotional expressions, and overall demeanor. Respond with ’Yes’ if (illness) is indicated, otherwise respond with ’No’. Respond with one word only! "(input)" Prompt 3 (P3) As a psychologist, analyze the provided interview (text/audio): (input) Does the interviewee exhibit clear symptoms of (illness)? Answer with ’Yes’ or ’No’ only without explaining your reasoning; use only one word!
@@ 4.6.2 Severity Prompts
In the following section, Figure 5 presents the PHQ_Score Mapped (Depression severity) Prompts, which are used to assess the severity of depression in participants. These prompts ask the model to categorize the level of depression based on the input provided, whether in text or audio format, using the PHQ-8 scoring system. The model classifies the participant’s depression into one of five categories: minimal or no depression (0), mild depression (1), moderate depression (2), moderately severe depression (3), and severe depression (4). These categories provide a nuanced view of the participant’s mental health by determining the extent of their depressive symptoms. In figures 5 and 6, the term "input" refers to either a transcription of the participant’s spoken responses or the original audio files used in the analysis.
Prompt 1 (P1) The following (input) is from an interaction between an interviewer and an interviewee. These interviews were collected as part of a larger effort to create a computer agent that interviews people and identifies verbal and nonverbal indicators of mental illnesses. The (input) is: ’{input}’. Based on the (input), classify the severity of depression into: 0 (Minimal or no depression): No or very few signs. 1 (Mild depression): Mostly mild symptoms. 2 (Moderate depression): Moderate symptoms impacting daily life. 3 (Moderately severe depression): Noticeable symptoms hindering daily life. 4 (Severe depression): Severe, pervasive symptoms greatly impairing function. Answer with one number only! Prompt 2 (P2) Analyze the following interview (input) for symptoms of depression. Focus on language patterns, emotional expressions, and demeanor. The (input) is: ’{input}’. Classify severity into: 0 (Minimal or no depression): No or very few signs. 1 (Mild depression): Mostly mild symptoms. 2 (Moderate depression): Moderate symptoms impacting daily life. 3 (Moderately severe depression): Noticeable symptoms hindering daily life. 4 (Severe depression): Severe, pervasive symptoms greatly impairing function. Answer with one number only!
@@
Prompt 1 (P1) The following (input) is from an interaction between an interviewer and an interviewee. These interviews were collected as part of a larger effort to create a computer agent that interviews people and identifies verbal and nonverbal indicators of mental illnesses. The (input) is: ’{input}’. Based on the (input), classify the severity of PTSD into: 0 (Little to no severity): No significant signs of PTSD symptoms; little to no impact on daily life. 1 (Moderately severe): Moderate symptoms of PTSD with a noticeable impact on behavior and daily routines. 2 (High severity): Severe PTSD symptoms with significant impact, disrupting daily life and possibly requiring intervention. Answer with one number only, 0 to 2, corresponding to these categories. Prompt 2 (P2) Analyze the following interview (input) to determine if the interviewee exhibits symptoms of PTSD. Focus on language patterns, emotional expressions, and overall demeanor. The (input) is: ’{input}’. Based on the (input), classify the severity of PTSD into: 0 (Little to no severity): No significant signs of PTSD symptoms; little to no impact on daily life. 1 (Moderately severe): Moderate symptoms of PTSD with a noticeable impact on behavior and daily routines. 2 (High severity): Severe PTSD symptoms with significant impact, disrupting daily life and possibly requiring intervention. Answer with one number only, 0 to 2, corresponding to these categories.
Figure 6: PTSD severity classification prompts.
@@ 4.6.3 Multiclass Prompts
Prompt 1 (P1) The following (input) is from an interaction between an interviewer and an interviewee. These interviews were collected as part of a larger effort to create a computer agent that interviews people and identifies verbal and nonverbal indicators of mental illnesses. The (input) is: ’{input}’. The interviewee could have multiple illnesses at the same time. Analyze the transcript and identify if the interviewee has Depression, PTSD, both, or is Normal. Respond with only one of these words: "Depressed", "PTSD", "Normal", "Depressed and PTSD".
Prompt 2 (P2) Analyze the following interview (input) to determine if the interviewee exhibits clear symptoms of Depression or PTSD according to the provided guidelines. Guidelines: The interviewee may have multiple illnesses at the same time or be normal. Concise Response: Respond with one of these four words only: "Depressed", "PTSD", "Depressed and PTSD", "Normal". No Explanations: Do not provide explanations for your assessment. Ambiguity: If the post is unclear, choose the most probable label. The (input) of the interview: ’{input}’. Only answer with one of the specified words.
@@ 4.6.4 Few-Shot Prompts
  • Severity and Multi-label Tasks: We strategically selected samples that were near-miss classifications by the model during zero-shot evaluations. This selection was based on samples where the model’s predictions were off by just one label, indicating a subtle misunderstanding of the distinguishing features between closely related classes. This methodological choice aimed to challenge the models with complex examples where minor nuances in features are decisive, thereby enhancing the robustness and precision of the models through few-shot learning.
@@
Example Few-shot Prompt (Prompt 1 for Binary Depression Detection) The following transcript is from an interaction between an interviewer and an interviewee. These interviews were collected as part of a larger effort to create a computer agent that interviews people and identifies verbal and nonverbal indicators of mental illnesses. If the text appears to be for a person who has Depression, answer with ’Yes’; if not, answer with ’No’. Only answer with Yes or No; respond with one word only! Here are 3 samples from these interviews and their labels. Use them as a reference: First sample transcription: (sample transcription) First sample label: No Second sample transcription: (sample transcription) Second sample label: Yes Third sample transcription: (sample transcription) Third sample label: Yes Label the following transcription: ’sample to be labeled’.
@@ 4.7 Parsers
Among the models used in this study, the larger models like Llama 3.1 70 Band Phi-3.5-MoE performed well when following task guidelines, typically adhering to the expected response formats. Conversely, some smaller models, such as Gemma 2 9B and Phi-3.5-Mini, struggled to maintain this level of compliance. These models frequently provided additional explanations, deviating from the expected outputs. Smaller models were less consistent, requiring additional handling to extract the necessary information.
@@
  • Binary Detection: The parser was designed to search the LLM outputs for the text "yes" or "no." If one of these options appeared in the response, it was taken as the final output. If both "yes" and "no" seemed simultaneously, or if neither were found, the model output was deemed invalid.
  • Severity Detection: For this task, the parser focused on extracting numerical values that corresponded to the specified severity range. If a single valid number was detected, it was accepted as the answer. However, if multiple numbers appeared or if the number fell outside of the allowed range, the output was flagged as invalid.
@@ 4.8 Evaluation Metrics
To evaluate the performance of the Large Language Models (LLMs) in tasks with uneven class distributions, as detailed in the dataset section (see Section 3), we prioritize Balanced Accuracy (BA) as our primary evaluation metric. This measure, calculated as the average recall obtained across each class, fairly reflects the model’s effectiveness in identifying both prevalent and rare conditions. By employing Balanced Accuracy, we ensure a comprehensive evaluation of the LLMs’ capabilities.
@@
Here, $i$ represents an index for each class, ranging from 0 to N1N-1, where $N$ is the total number of classes. The recall for each class is computed to assess the model’s ability to correctly identify samples of that particular class.
@@
For binary classification tasks, we additionally use the F1 Score to evaluate model performance. The F1 Score is crucial as it provides a balance between precision and recall, making it a valuable metric for situations where the cost of false positives and false negatives is high. This metric is critical in mental health assessments, where accurately distinguishing between conditions such as depression and PTSD is critical.
@@
$\text{{F1 Score}}=2\times\left(\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}\right)$
While weighted F1 score is defined as
$\text{{Weighted F1 Score}}=\sum_{i=1}^{N}w_{i}\times\text{F1 Score}_{i}$
where $N$ is the number of classes, $w_{i}$ is the weight for class $i$, and $\text{F1 Score}_{i}$ is the F1 score for class $i$. The weight for each class, $w_{i}$, is defined as:
@@
To comprehensively assess the performance differences between each modality in our analysis, we employed two key metrics: the Modal Superiority Score (MSS) and the Disagreement Resolution Score (DRS). These metrics are instrumental in quantifying the relative efficacy of individual and combined modalities in making correct predictions, particularly in the face of disagreement, by applying these metrics, which are detailed in the equations below, we gain valuable insights into which modalities perform better or worse and explore whether combining modalities enhances the predictive accuracy and giving insights on how a modality improves upon another.
Modal Superiority Score (MSS): The Modal Superiority Score (MSS) quantifies the net superiority of one modality over another by comparing how often each modality correctly predicts outcomes when the other does not. This metric can be applied not only to comparisons between individual modalities, such as Audio versus Text, but also in evaluating the performance of a combined modality (e.g., Audio+Text) against individual modalities and the collective agreement of these modalities (e.g., cases where both Audio and Text are either correct or incorrect). This comprehensive application of MSS allows for assessing the relative strength of combined modalities over both their constituent individual modalities and their collective concordance. MSS values can be positive or negative, with a positive value indicating that modality A performs better than modality B, and a negative value suggesting the opposite.
[…]
@@
$\text{DRS}=\left(\frac{\text{Correctly Resolved}-\text{Incorrectly Resolved}}{\text{Total Number of Disagreements}}\right)\times 100\%$
@@ 5 Results & Discussion
In the upcoming sections, we analyze the performance of various models for several mental health classification tasks, including Binary Depression, Binary PTSD, Depression Severity, and PTSD Severity. The aim is to determine which models perform best under different conditions across both text and audio modalities, with no preprocessing or fine-tuning applied, relying solely on zero-shot inference.
@@ 5.1 Binary Depression Classification Results
Table 6: Report on model performances for Depression binary classification across text and audio modalities and their combination. Details about the specific prompts used can be found in Figure 4. The underlined bold values are the best score for the specific column.
@@
Text
Model BA F1 BA F1 BA F1
Llama 3.1 70B 71.9% 0.61 74.7% 0.64 65.7% 0.54
@@
Deepseek_R1 79.8% 0.75 74.8% 0.68 77.5% 0.78
Gemini Flash Lite 76.9% 0.73 69.1% 0.58 77.3% 0.75
Gemini 2 Flash 74.2% 0.68 70.42% 0.61 74.2% 0.74
Gemini Flash Thinking 73.4% 0.64 70.7% 0.64 76.6% 0.74
Gemini 2.5 Pro 74.8% 0.68 73% 0.65 77.3% 0.77
Llama 3.3 70B 72.9% 0.66 74.1% 0.66 76.3% 0.73
Audio
Gemini 1.5 Pro 76.1% 0.66 74.5% 0.64 74% 0.64
@@
Gemini Flash Lite 72.7% 0.62 75.2% 0.65 67.6% 0.56
Gemini 2 Flash 72.5% 0.67 70.8% 0.63 76.8% 0.74
Gemini Flash Thinking 74.5% 0.73 75.3% 0.74 67.5% 0.69
Gemini 2.5 Pro 75.5% 0.70 69.3% 0.60 78.9% 0.76
Audio and Text
Gemini 1.5 Pro 74% 0.64 77% 0.67 77.3% 0.67
@@
Gemini 2 Flash 73.5% 0.63 69.5% 0.60 75.3% 0.66
Gemini Flash Thinking 74.4% 0.64 69.5% 0.60 78% 0.69
Gemini 2.5 Pro 77.8% 0.74 72.8% 0.65 78.9% 0.77
Table 6 presents the results for binary depression classification across Text, Audio, and combined Audio+Text modalities, evaluated using Balanced Accuracy (BA) and F1-score. The results span three prompt variants (detailed in Figure 4), with the best-performing scores for each column highlighted in bold and underlined.
The highest overall Balanced Accuracy (BA) of 79.8% was achieved by Deepseek_R1 using the Text modality with Prompt 1. In terms of F1-score, the best result of 0.78 was also obtained by the same model using Prompt 3. The Audio modality, while slightly behind Text in maximum BA, demonstrated strong capability with Gemini 2.5 Pro, reaching 78.9% BA and 0.76 F1 with Prompt 3. Notably, combining Text and Audio resulted in competitive results, with Gemini 2.5 Pro attaining 78.9% BA and 0.77 F1 using Prompt 3, highlighting the strength of multimodal integration for this task.
Deepseek_R1 consistently delivered the highest scores, particularly excelling with Prompts 1 and 3. Gemini 2.5 Pro exhibited strong, balanced performance across all modalities, notably excelling in Audio-only and multimodal setups. These models demonstrated consistent accuracy and high F1 scores, indicating their robustness in binary depression classification. Llama 3.3 70B and GPT-4o mini offered moderate performance, often showing competitive F1 scores but not achieving the peak BA observed in Deepseek_R1 and Gemini models. Phi-3.5-MoE mini and older Gemini Flash versions showed more variability and generally lower consistency across prompts and modalities. The latest Pro versions in the Gemini family, especially 2.5 Pro, outperformed earlier variants (e.g, Flash 1.5 and Thinking) across modalities, indicating notable architectural or optimization improvements.
Thinking models, such as Gemini Flash Thinking, presented mixed results. Although they occasionally performed well (e.g., 76.6% BA in Text, Prompt 3), they generally did not outperform non-thinking counterparts. Indicates that incorporating deliberative processing may not always yield improvements in structured clinical classification tasks, suggesting that the utility of "Thinking" variants is context-dependent.
To evaluate the impact of prompt choice, we calculated the average performance for each metric across Prompt 1, Prompt 2, and Prompt 3. The results indicate that Prompt 3 showed a slight overall advantage in F1 scores, while Prompt 1 was more reliable in achieving the highest balanced accuracy. This pattern suggests that while Prompt 3 is particularly effective in fine-tuning F1-based classification, Prompt 1 remains superior when prioritizing balanced accuracy.
This evaluation affirms the effectiveness of large-scale language and multimodal models in depression classification tasks. Text remains the most potent single modality; however, combining it with Audio data enhances performance, especially with models like Gemini 2.5 Pro. Prompt variation plays a critical role, and newer model architectures substantially outperform earlier iterations. These findings highlight the importance of model selection, prompt tuning, and modality fusion in advancing affective computing tasks.